An old bus and a new bus left the same station heading in the same direction at 5:00 am. The old bus was traveling at 40 miles per hour, and the new bus was traveling at 50 miles per hour. If the busses keep going at those rates without stopping, how far apart will they be at 10:00 am?
Answer by stanbon(75887)

An old bus and a new bus left the same station heading in the same direction at 5:00 am. The old bus was traveling at 40 miles per hour, and the new bus was traveling at 50 miles per hour. If the busses keep going at those rates without stopping, how far apart will they be at 10:00 am?
---------------
Old bus distance:
dist = rate*time
d = 40*5 = 200 miles
---------------------
New bus distance:
dist = rate*time
d = 50*5 = 250 miles
---------------------
Distance apart at 10AM: 250-200 = 50 miles
=============================
Cheers,
Stan H.